A body falls for 5s from rest. If the acceleration due to gravity of earth ceases to act, the distance it travels in the next 3s is
73.5m
294m
147m
49m

Grade:11

1 Answers

Neeti
8 years ago
Initially the body starts it’s motion from rest therefore u = 0.
let this journey be A, when accelaration due to earth is acting on the body.
time = 5seconds. Let’s find v
v= u+ at
= 0 + 9.8 x 5
=49m/s
 
when accelaration stops acting, the velocity of the body is 49m/s
now the distance travelled by the body in the next 3s is
 
s= ut + 1/2gt2
  = 49 x 3 + ½ x 0 x t2
   = 49 x 3
 
147m
 
If you like the answer please approve it and feel free to ask any follow-up question :)